# Learning holographic QCD with unflavoured meson spectra

### v1

@@ 2.1.2 Vector and Axial-Vector Mesons
$\partial_{z}\left(e^{-B_{1}(z)}\partial_{z}V_{n}(z)\right)+m_{Vn}^{2}e^{-B_{1}(z)}V_{n}(z)=0,$ (14)
where $B_{1}(z)=\phi(z)-A(z)$, and $m_{Vn}^{2}$ are the mass eigenvalues corresponding to the masses of the vector meson. To bring this equation to a Schrödinger-like form we substitute, $V_{n}(z)=e^{B_{1}(z)/2}\psi_{Vn}(z)$ and obtain,
$\partial_{z}^{2}\psi_{Vn}(z)+V_{\rho}\psi_{Vn}(z)=m_{Vn}^{2}\psi_{Vn}(z),$ (15)
@@
The mass of the vector mesons are computed by demanding Dirichlet boundary condition for $\psi_{Vn}$ at UV and for large $z$.
@@
$\partial_{z}(e^{-B_{1}(z)}\partial_{z}a_{n})+\left(m_{an}^{2}-g_{5}^{2}e^{2A(z)}v^{2}(z)\right)e^{-B_{1}(z)}a_{n}=0.$ (17)
Here we denote the axial-vector meson masses by $m_{an}^{2}$. Substituting $a_{n}(z)=e^{B_{1}(z)/2}\psi_{An}(z)$, we obtain,
$\partial_{z}^{2}\psi_{An}(z)+V_{a_{1}}\psi_{An}(z)=m_{An}^{2}\psi_{An}(z),$ (18)
such that $\psi_{An}(z)$ satisfies Dirichlet boundary condition at UV and at large $z$. The effective potential is given by,
@@ 2.1.3 Higher-Spin Mesons
$\partial_{z}\left(e^{-B_{G}(z)}\partial_{z}G_{n}(z)\right)+m_{Gn}^{2}e^{-B_{G}(z)}G_{n}(z)=0,$ (20)
where $B_{G}(z)=\phi(z)-(2G-1)A(z)$. For a spin-2 meson like the $a_{2}$ meson, we have $G=2$ and substituting $G_{n}(z)=e^{B_{G}(z)/2}\psi_{Gn}(z)$, we obtain the Schrödinger-like equation,
$\partial_{z}^{2}\psi_{Gn}(z)+V_{a_{2}}\psi_{Gn}(z)=m_{Gn}^{2}\psi_{Gn}(z)\ ,$ (21)
@@
with $B_{2}(z)=\phi(z)-3A(z)$. The mass eigenvalues $m_{Gn}^{2}$ of the spin-2 meson are obtained again by solving the equation of motion and imposing the Dirichlet boundary conditions on the wave functions.
With these, in the next section we discuss a neural network architecture to solve the Schrödinger-like equations, along with the boundary conditions, to obtain the warp factor, dilaton profile, and the chiral symmetry breaking potential by training the model with the masses $m_{Sn}$, $m_{Vn}$, $m_{An}$ and $m_{Gn}$ in Table 1. Moreover, the network will predicted the mass spectrum of pions $m_{\pi n}$ and match with the experimental valeus.
@@ 3.1 Model Architecture
This transformation ensures the correct boundary condition at the UV region. The denominator in $v(z)$ is introduced in order to prevent numerical instability in the IR region.
@@
where, $\sigma$ is the sigmoid function and $a=0.01$ GeV3 and $b=0.03$ GeV3. The quark mass $m_{q}$ is then computed using the GMOR relation given in Eq. 7 and this is, in turn, used to compute the $\alpha$ and $\beta$ values from Eq. 6.
For a background $A(z)$ and the VEV $v(z)$ trained with Eq.23, the $B_{2}(z)$ could be computed from Eq. 4 as,
@@
The coefficients $k_{1}$ and $k_{2}$ are the trainable parameters. Although we trained the model using only the quartic interaction term, it was found that adding a cubic interaction term significantly reduced the loss. This may indicate either insufficient hyperparameter tuning in the quartic-only case or that a cubic interaction is genuinely required in the holographic model.
@@
$L=\log(1+l^{2})\ ,$ (28)
to ensures that it remains positive while also preventing it from taking excessively large values, while training.
@@ 3.2 Loss Functions
where $\omega_{n}$ are the experimental values of the meson masses as given in [47]. The Loss function now becomes,
@@
where $N_{\mu}$ is the number of excited mesons considered for each kind, and $\Lambda_{\mu,n}$ is a hyperparameter that controls the weightage of each meson mass in the total loss function. We have employed an adaptive weighting scheme where $\Lambda_{\mu,n}$ at each training step is proportional to the loss contributed by that particular mass (See Appendix B: Eq. 49). This ensures that the model focuses more on the masses that are harder to fit. The total mass loss function used in the training is given by
@@
To confine the wavefunction within the AdS space and ensure normalizability of the wavefunction at large $z$, we expect the potentials $V_{\mu}(z)$ to be non-decreasing functions of $z$ in the IR region. This can be imposed by demanding that $\partial_{z}V_{\rho}(z)$ and $\partial_{z}V_{a_{2}}(z)$ are positive at large $z$. This automatically also ensures that the potentials given in Eq. 10 and Eq. 19 and are also confining. To impose this condition, we add another penalty term to the loss function given by
@@
where $z$ is chosen randomly at each training step at a sufficiently large value close to $z_{max}$. We expect $A(z)$ to be negative and decreasing at the IR from warped geometry. This is imposed by adding another penalty term,
$\mathcal{L}_{A}=\text{ReLU}\left(A(z_{max})+\ln(z_{max})-\Delta\right)\ ,$ (35)
where $\Delta$ is a constant. The total loss function now becomes,
@@
where $\Lambda_{v^{\prime}}$, $\Lambda_{\text{pot}}$, and $\Lambda_{A}$ are hyperparameters that control the weightage of the penalty terms. The model is trained to minimize the total loss function $\mathcal{L}_{\text{tot}}$ by optimizing the neural network’s weights and biases, as well as the trainable parameters $\theta$, $_{1}, kk$, and $L$.
@@ 4 Numerical Results
We perform a hyperparameter search for the initial values of $k_{1}$ and $k_{2}$ using Optuna’s built-in TPE sampler, in the range [15,15][-15,15] (See Appendix B). We observe that the model converges to a lower loss value for negative $k_{1}$ and positive $k_{2}$, resulting in a bounded scalar field potential. The best initial guess was found to be around $k_{1}=-7.4$ and $k_{2}=12.2$. With this initial guess, we train our model several times and select only those runs in which the total loss converges to a value below $0.01$. We have selected 10 such runs, whose training loss and the trained model parameters ($k_{1}$, $k_{2}$, $L$, $\theta$) corresponding to each run are listed in Table 5. Further details of the training procedure, along with the hyperparameters used, are provided in Appendix B.
Figure 1: The learned functions $v(z)$ (left) and $A(z)$ (right) after training. The shaded regions represent the standard deviation over different runs, and the solid lines represent the mean value.
$L$ $1.7051\pm 0.2638\ GeV^{-1}$ $\theta$ $2.4094\pm 1.3162$
$k_{1}$ $7.7745\pm 1.0499$ $k_{2}$ $17.0736\pm 2.7509$
Table 2: The trained model parameters $L=\log{(1+l^{2})}$, $\theta$, $k_{1}$ and $k_{2}$. The values are given as mean $\pm$ standard deviation over different runs.
@@
$\Sigma=(0.3004\pm 0.0127\ \text{GeV})^{3}\quad\text{and}\quad m_{q}=0.0031\pm 0.0005\ \text{GeV},$ (37)
where the values are given as mean $\pm$ standard deviation over different runs. With the given $v(z)$ and $A(z)$, we can obtain $d\phi/dz$ which is plotted in Fig. 2 (left). We observe a large variation in $d\phi/dz$ across different runs, especially in the IR region. Integrating $d\phi/dz$ gives us $\phi(z)$ up to an additive constant, also given in Fig. 2 (right). The $\phi(z)$, unlike $d\phi/dz$, shows a much more symmetric spread across different runs. We have plotted a linear and quadratic fit to the mean value of $\phi(z)$. Additionally, we plot the effective potentials for the $\rho$, $a_{1}$, $a_{2}$, and $f_{0}$ mesons in Fig. 3. Like the $d\phi/dz$, these potentials also exhibit a broader and more skewed variation across different runs.
Figure 2: The plot of $d\phi/dz$ (left) and $\phi(z)$ (right). The solid line represents the best fit corresponding to the minimum value of $\mathcal{L}_{\text{mass}}^{(\rho,a_{1},a_{2},f_{0})}$. Unlike Figs. 1 and 2, where the spread was symmetric and shown as mean $\pm$ standard deviation, the $d\phi/dz$ here exhibit a skewed variation, which is why we only plot the $d\phi/dz$ corresponding to the best run. However, the $\phi(z)$ (right panel) obtained shows a symmetric spread, which is represented by the mean. The dotted line and the dashed line in the right panel represent a linear and quadratic fit to the mean value of $\phi(z)$, respectively.
Figure 3: Effective potentials for the $\rho$, $a_{1}$, $a_{2}$, and $f_{0}$ mesons. The solid line represents the best fit corresponding to the minimum value of $\mathcal{L}_{\text{mass}}^{(\rho,a_{1},a_{2},f_{0})}$.
The predicted spectra of the trained masses of the $\rho$, $a_{1}$, $a_{2}$, and $f_{0}$ mesons obtained from our model are given in Table 3 and are plotted with respect to n (radial excitation number) in Fig. 4. The predicted masses show good agreement with the experimental values, with a maximum error of around 4.23%.
@@
$\rho$ 1 $0.79\pm 0.01$ 2.93%
2 $1.46\pm 0.01$ 0.83%
3 $1.61\pm 0.01$ 2.84%
4 $1.77\pm 0.02$ 4.23%
5 $1.91\pm 0.03$ 0.70%
6 $2.11\pm 0.02$ 1.91%
$a_{1}$ 1 $1.25\pm 0.01$ 1.06%
2 $1.64\pm 0.01$ 0.25%
$a_{2}$ 1 $1.31\pm 0.01$ 0.46%
2 $1.66\pm 0.01$ 2.20%
$f_{0}$ 1 $0.49\pm 0.00$ 1.36%
2 $98\pm 0.01 0.0$2%
3 $1.39\pm 0.03$ 1.75%
4 $1.50\pm 0.04$ 0.20%
5 $1.67\pm 0.02$ 1.96%
6 $1.83\pm 0.02$ 0.12%
7 $1.98\pm 0.03$ 1.24%
8 $2.13\pm 0.02$ 0.97%
9 $2.28\pm 0.02$ 1.01%
10 $2.42\pm 0.03$ 0.88%
Table 3: Predicted meson masses (in GeV). The predictions are reported as the mean $\pm$ standard deviation across different runs. The % Error is calculated with respect to the mean value of the predicted mass.
State Predicted Mass (GeV) Experimental
Mass (GeV)
Best Fit:
Lowest $\mathcal{L}_{\text{mass}}^{(\pi)}$ (GeV)
@@
To test our model, we predict the masses of the pseudo-scalar meson $\pi$. We use the finite difference method to solve the coupled second-order differential equation given in Eq. 13 using the learned parameters and functions. This time, we obtain a matrix for a generalized eigenvalue equation (see Appendix A for more details), which we solve using SciPy’s built-in function scipy.linalg.eigvals.
The predicted masses eigenvalues $m_{n}$ obtained (after converting to physical units) using our trained model are given in Table 4 and plotted in Fig. 5. Here, the best fit values are obtained for the run with the lowest value of $\mathcal{L}_{mass}^{(\pi)}$. The Loss corresponding to the $\pi$ meson masses for each run is given in Table 5. Although we have not trained our model to fit the $\pi$ meson masses, the predicted values match the experimental data fairly well, except for the first excited state, where the model has overshoot the mass by a large amount.
@@ 5 Conclusion
The bottom-up approach in holographic QCD explores the functional space of all possible background metrics and scalar potentials, searching for functions that best fit the QCD observables. With enough phenomenological and theoretical constraints, the bottom-up approach can severely restrict the allowed background functions and may select a class of viable models, though not necessarily a unique one. We have provided a deep-learning architecture that searches in the possible functional space of the warp-factor $A(z)$, the dilaton profile $\phi(x)$, along with the possible values of cubic ($k_{1}$) and quartic ($k_{2}$) terms in the scalar potential, and the quark condensate $\Sigma$. This space is constrained by fitting the mass spectrum of the unflavored mesons ($\rho,a_{1},a_{2}$ and $f_{0}$) and their excitations.
The duality between AdS/QCD and deep neural networks, as shown in [48], raises intriguing questions about their physical origins and suggests a new methodology for investigating both frameworks. This was pioneered in [27], where the authors trained for the masses of the $\rho$ and $a_{1}$ meson, through which the $A(z)$ and $\phi(z)$ were determined. They solved the differential equation of the KK mode using the Euler method, which, unfortunately, results in the accumulation of local truncation error at each iteration. The eigenvalues are calculated by traversing all the values of masses to search for the mass for which KK modes are normalizable. Since the axial sector was not studied, the information about the chiral symmetry breaking by the bulk scalar field is not well understood. Including the higher meson masses and axial sector, along with the $f_{0}$ mesons, becomes computationally challenging because the network’s structure is determined by the theory itself.
In contrast, we propose here to train the architecture using the finite difference method to directly obtain the mass eigenvalue, making it more efficient, less error-prone, and more extensible to higher excitation states of the meson tower. Additionally, the introduction of the scalar vev has enabled us to determine the masses of all unflavored mesons. This strategy is flexible and highly tunable, which leaves scope for employing additional constraints on the functional space in the form of additional loss functions.
@@
Though the $z^{2}$ behavior for the dilaton profile is anticipated by several phenomenological models [12, 17, 18, 23] for linear confinement, string theoretic models often assume a linear profile. In this work, the predicted dilaton, in Fig.2, exhibits a profile in between the linear and quadratic types, which is positive throughout the region, making it consistent with the null energy condition [39]. Moreover, it predicted the pion mass and its excitations with good accuracy. As mentioned earlier, our numerical experiment suggests that for a more precise fitting of the mass spectra, a cubic interaction, in addition to the quartic interaction, is necessary.

### v2

@@ 2.1.2 Vector and Axial-Vector Mesons
$\partial_{z}\left(e^{-B_{1}(z)}\partial_{z}V_{n}(z)\right)+m_{\rho,n}^{2}e^{-B_{1}(z)}V_{n}(z)=0,$ (14)
where $B_{1}(z)=\phi(z)-A(z)$, and $m_{\rho,n}^{2}$ are the mass eigenvalues corresponding to the masses of the vector meson. To bring this equation to a Schrödinger-like form we substitute, $V_{n}(z)=e^{B_{1}(z)/2}\psi_{\rho,n}(z)$ and obtain,
$\partial_{z}^{2}\psi_{\rho,n}(z)+V_{\rho}\psi_{\rho,n}(z)=m_{\rho,n}^{2}\psi_{\rho,n}(z),$ (15)
@@
The masses of the vector meson are computed by demanding the Dirichlet boundary condition $\psi_{\rho,n}\rightarrow 0$ at UV and for large $z$.
@@
$\partial_{z}(e^{-B_{1}(z)}\partial_{z}a_{n})+\left(m_{a_{1},n}^{2}-g_{5}^{2}e^{2A(z)}v^{2}(z)\right)e^{-B_{1}(z)}a_{n}=0.$ (17)
Here we denote the axial-vector meson masses by $m_{a_{1},n}^{2}$. Substituting $a_{n}(z)=e^{B_{1}(z)/2}\psi_{a_{1},n}(z)$, we obtain,
$\partial_{z}^{2}\psi_{a_{1},n}(z)+V_{a_{1}}\psi_{a_{1},n}(z)=m_{a_{1},n}^{2}\psi_{a_{1},n}(z),$ (18)
such that $\psi_{a_{1},n}(z)$ satisfies Dirichlet boundary condition at UV and at large $z$. The effective potential is given by,
@@ 2.1.3 Higher-Spin Mesons
$\partial_{z}\left(e^{-B_{G}(z)}\partial_{z}G_{n}(z)\right)+m_{a_{G},n}^{2}e^{-B_{G}(z)}G_{n}(z)=0,$ (20)
where $B_{G}(z)=\phi(z)-(2G-1)A(z)$. For a spin-2 meson like the $a_{2}$ meson, we have $G=2$ and substituting $G_{n}(z)=e^{B_{G}(z)/2}\psi_{a_{2},n}(z)$, we obtain the Schrödinger-like equation,
$\partial_{z}^{2}\psi_{a_{2},n}(z)+V_{a_{2}}\psi_{a_{2},n}(z)=m_{a_{2},n}^{2}\psi_{a_{2},n}(z)\ ,$ (21)
@@
with $B_{2}(z)=\phi(z)-3A(z)$. The mass eigenvalues $m_{a_{2},n}^{2}$ of the spin-2 meson are obtained by solving the equation of motion and imposing the Dirichlet boundary conditions on the wave functions.
With these, in the next section we discuss a neural network architecture to solve the Schrödinger-like equations, along with the boundary conditions, to obtain the warp factor, dilaton profile, and the chiral symmetry breaking potential by training the model with the masses $m_{f_{0},n}$, $m_{\rho,n}$, $m_{a_{1},n}$ and $m_{a_{2},n}$ in Table 1. Moreover, the network will be used to predict the pion mass spectrum, $m_{\pi,n}$, and compare them with experimental values.
@@ 3.1 Model Architecture
This transformation ensures the correct boundary condition at the UV region. The denominator in $v(z)$ is introduced in order to prevent numerical instability in the IR region. This does not affect the UV behavior of v(z) up-to order of $\mathcal{O}(z^{4})$.
@@
where, $\sigma$ is the sigmoid function and $a=0.01$ GeV3 and $b=0.03$ GeV3. The quark mass $m_{q}$ is then computed using the GMOR relation given in Eq. 7, and this is, in turn, used to compute the $\alpha$ and $\beta$ values from Eq. 6.
For a given background $A(z)$ and the VEV $v(z)$, the $B_{2}(z)$ could be computed from Eq. 4 as,
@@
The coefficients $k_{1}$ and $k_{2}$ are the trainable parameters. We have observed that adding a cubic interaction term along with a quartic term significantly reduces the loss. This may indicate either insufficient hyperparameter tuning in the quartic-only case or that a cubic interaction is genuinely required in the holographic model.
@@
$L=\log(e+l^{2})\ ,$ (28)
to ensure that its value remains greater than 1, while also preventing it from taking excessively large values, while training.
@@ 3.2 Loss Functions
where $\omega_{n}$ are the experimental values of the meson masses as given in [54]. The Loss function is then defined as,
@@
where $N_{\mu}$ is the number of excited mesons considered for each kind, and $\Lambda_{\mu,n}$ is a hyperparameter that controls the weightage of each meson mass in the total loss function. We have employed an adaptive weighting scheme where $\Lambda_{\mu,n}$ at each training step is proportional to the loss contributed by that particular mass (See Appendix B: Eq. 49). This ensures that the model focuses more on the masses that are harder to fit. Further, we have scaled the error for the first mass of the $f_{0}(500)$ meson $E_{f_{0},1}$ by a factor of 0.1, as it is an outlier in the mass spectrum and the experimental value is not well determined. The total mass loss function used in the training is given by
@@
To confine the wavefunction within the AdS space and ensure normalizability of the wavefunction at large $z$, we expect the potentials $V_{\mu}(z)$ to be non-decreasing functions of $z$ in the IR region. This can be imposed by demanding that $\partial_{z}V_{\rho}(z)$ and $\partial_{z}V_{a_{2}}(z)$ are positive at large $z$. This automatically also ensures that the potentials given in Eq. 10 and Eq. 19 are also confining. To impose this condition, we add another penalty term to the loss function given by
@@
where $z$ is chosen randomly at each training step at a sufficiently large value close to $z_{max}$. We expect $A(z)$ to be negative and decreasing at the IR scale due to warped geometry. This is imposed by adding another penalty term,
$\mathcal{L}_{A}=\text{ReLU}\left(A(z_{max})+\ln(z_{max})-\Delta\right)+\text{ReLU}\left(\partial_{z}A(z_{max})\right),$ (35)
where $\Delta$ is a constant we choose to be 1.5. The total loss function now becomes,
@@
where $\Lambda_{v^{\prime}}$, $\Lambda_{\text{pot}}$, and $\Lambda_{A}$ are hyperparameters that control the weightage of the penalty terms. The model is trained to minimize the total loss function $\mathcal{L}_{\text{tot}}$ by optimizing the neural network’s weights and biases, as well as the trainable parameters $\theta$, $_{1}, kk$, and $L$. A schematic diagram summarizing the entire architecture is shown in Figure 1.
Figure 1: Summary of the Neural Network training architecture.
@@ 4 Numerical Results
We perform a hyperparameter search for the initial values of $k_{1}$ and $k_{2}$ using Optuna’s built-in TPE sampler, in the range [20,20][-20,20] (See Appendix B). We observe that the model converges to a lower loss value for negative $k_{1}$ and positive $k_{2}$, resulting in a bounded scalar field potential. The best initial guess was found to be around $k_{1}=-2.2$ and $k_{2}=12.0$. With this initial guess, we train our model several times and select only those runs in which the total loss converges to a value below $0.01$. We have selected 10 such runs, whose training loss and the trained model parameters ($k_{1}$, $k_{2}$, $L$, $\theta$) corresponding to each run are listed in Table 5. Further details of the training procedure, along with the hyperparameters used, are provided in Appendix B.
Figure 2: The learned functions $v(z)$ (left) and $A(z)$ (right) after training corresponding to the best fit to the $\pi$ meson masses.
$L$ $1.234\pm 0.229\ GeV^{-1}$ $\theta$ $3.697\pm 1.305$
$k_{1}$ $4.332\pm 1.022$ $k_{2}$ $8.600\pm 2.355$
Table 2: The trained model parameters $L=\log{(e+l^{2})}$, $\theta$, $k_{1}$ and $k_{2}$. The values are given as mean $\pm$ standard deviation over different runs.
@@
$\Sigma=(0.3071\pm 0.0042\ \text{GeV})^{3}\quad\text{and}\quad m_{q}=0.0029\pm 0.0001\ \text{GeV},$ (37)
where the values are given as mean $\pm$ standard deviation over different runs. With the given $v(z)$ and $A(z)$, we can obtain $d\phi/dz$ which is plotted in Fig. 3 (left). Integrating $d\phi/dz$ gives us $\phi(z)$ up to an additive constant, also given in Fig. 3 (right). We have plotted a linear and quadratic fit to the function $\phi(z)$. Additionally, we plot the effective potentials for the $\rho$, $a_{1}$, $a_{2}$, and $f_{0}$ mesons in Fig. 4.
Figure 3: The plot of $d\phi/dz$ (left) and $\phi(z)$ (right) for the trained models corresponding to the best fit to the $\pi$ meson masses. The dotted line and the dashed line in the right panel represent a linear and quadratic fit to the mean value of $\phi(z)$, respectively.
Figure 4: Effective potentials (log scaled) for the $\rho$, $a_{1}$, $a_{2}$, and $f_{0}$ mesons corresponding to the best fit to the $\pi$ meson masses.
The predicted spectra of the trained masses of the $\rho$, $a_{1}$, $a_{2}$, and $f_{0}$ mesons obtained from our model are given in Table 3 and are plotted with respect to n (radial excitation number) in Fig. 5. The predicted masses agree well with the experimental values. The comparatively larger error for the $f_{0}(500)$ is not unexpected, as it reflects the larger experimental uncertainty of this broad resonance (400 to 500 MeV), and the reduced weightage assigned to its loss during training, which allows for greater variation in its predicted mass. Using the trained model, we predict the masses $f_{0}(1370)$, which we skipped in training due to the large uncertainty in its experimental value. The predicted mass of $f_{0}(1370)$ is $1.49\pm 0.02$ GeV, which is consistent with the given range of $1.2$ to $1.5$ GeV in the Summary Table of the Particle Data Group [54]. We further predict the higher excited states of the $\rho$ and $f_{0}$ mesons, which are not well established experimentally. The predicted masses of the $\rho(1900)$, $\rho(2150)$, $f_{0}(2200)$, $f_{0}(2330)$ and $f_{0}(2470)$ are given in Table 3 and are within the expected range of experimental uncertainties. We have omitted the state $\rho(1570)$ in the table, because it may be an OZI-violating decay mode of $\rho(1700)$.
@@
$\rho$ 11 $0.78\pm 0.01$ 0.95%
22 $1.46\pm 0.01$ 0.15%
33 $1.71\pm 0.01$ 0.37%
$4^{\prime}$ $1.93\pm 0.03$ 1.04%
$5^{\prime}$ $2.09\pm 0.01$ 0.41%
$a_{1}$ 11 $1.23\pm 0.01$ 0.20%
22 $1.67\pm 0.01$ 0.88%
$a_{2}$ 11 $1.31\pm 0.00$ 0.49%
22 $1.70\pm 0.01$ 0.33%
$f_{0}$ 11 $0.38\pm 0.03$ 19.18%
22 $0.98\pm 0.01$ 0.74%
$3^{\prime}$ $1.49\pm 0.02$ 1.54%
44 $1.54\pm 0.01$ 1.07%
55 $1.75\pm 0.01$ 0.87%
66 $1.95\pm 0.01$ 1.62%
$7^{\prime}$ $2.13\pm 0.03$ 2.40%
$8^{\prime}$ $2.34\pm 0.04$ 0.01%
$9^{\prime}$ $2.54\pm 0.08$ 2.77%
Table 3: Predicted meson masses (in GeV). The predictions are reported as the mean $\pm$ standard deviation across different runs. The % Error is calculated with respect to the mean value of the predicted mass. The unprimed states were used for model training.
State Predicted Mass (GeV)
(Mean $\pm$ Std)
Predicted Mass (GeV)
(Best fit: Lowest $\mathcal{L}_{\text{mass}}^{(\pi)}$)
@@
To validate our model, we predict the masses of the pseudo-scalar meson $\pi$. We use the finite difference method to solve the coupled second-order differential equation given in Eq. 13 using the learned parameters and functions. This time, we obtain a matrix for a generalized eigenvalue equation (see Appendix A for more details), which we solve using SciPy’s built-in function scipy.linalg.eigvals.
The predicted masses eigenvalues $m_{n}$ obtained (after converting to physical units) using our trained model are given in Table 4 and plotted in Fig. 6. Here, the best fit values are obtained for the run with the lowest value of $\mathcal{L}_{mass}^{(\pi)}$. The Loss corresponding to the $\pi$ meson masses for each run is given in Table 5. Although the $\pi$ meson masses are not included in the training loss, they are taken into account during the hyperparameter optimization procedure. The predicted values nevertheless show good agreement with experimental data, except for the first excited state, where the model significantly overshoots the mass.
@@ 5 Conclusion
The bottom-up approach in holographic QCD explores the functional space of all possible background metrics and scalar potentials, searching for functions that best fit the QCD observables. With enough phenomenological and theoretical constraints, the bottom-up approach can severely restrict the allowed background functions and may select a class of viable models. We have provided a deep-learning architecture that searches in the possible functional space of the warp-factor $A(z)$, the dilaton profile $\phi(x)$, along with the possible values of cubic ($k_{1}$) and quartic ($k_{2}$) terms in the scalar potential, and the quark condensate $\Sigma$. This space is constrained by fitting the mass spectrum of the unflavored mesons ($\rho,a_{1},a_{2}$ and $f_{0}$) and their excitations.
The duality between AdS/QCD and deep neural networks, as shown in [34], raises intriguing questions about their physical origins and suggests a new methodology for investigating both frameworks. This was pioneered in [7], where the authors trained for the masses of the $\rho$ and $a_{1}$ meson, through which the $A(z)$ and $\phi(z)$ were determined. They solved the differential equation of the KK mode using a discretized Euler-type scheme, which may accumulate numerical errors over long evolutions. The eigenvalues are calculated by traversing all the values of masses to search for the mass for which KK modes are normalizable. Since the axial sector was not studied, the information about the chiral symmetry breaking by the bulk scalar field is not well understood. Including higher-meson masses and the axial sector, along with the $f_{0}$ mesons, becomes computationally challenging because the network’s structure is determined by the theory itself.
In contrast, we propose here to train the architecture using the finite-difference method to directly obtain the mass eigenvalue, thereby making it more efficient, less error-prone, and more extensible to higher excitation states of the meson tower. Additionally, the introduction of the scalar vev has enabled us to determine the masses of all unflavored mesons. This strategy is flexible and highly tunable, leaving scope to impose additional constraints on the functional space via additional loss functions.
@@
Though the $z^{2}$ behavior for the dilaton profile is anticipated by several phenomenological models [43, 27, 67, 52] for linear confinement, string theoretic models often assume a linear profile. In this work, the predicted dilaton, in Fig. 3, exhibits a profile that increases more steeply than the quadratic behavior, and is positive throughout the region, making it consistent with the null energy condition [47]. Moreover, it predicted the pion mass and its excitations to a reasonable level of accuracy. As mentioned earlier, our numerical experiment suggests that for a more precise fitting of the mass spectra, a cubic interaction, in addition to the quartic interaction, is necessary.